Let W={(X,Y,Z)|x+y+z=0}

A, give a basis for W

B, what is the dimension of W



1
Expert's answer
2021-11-01T19:01:07-0400

Part A


"x+y+z=0\>"

"\implies\>x=-y-z"


"\begin{pmatrix}
 -y&-z \\
 y \\
z
\end{pmatrix}" ="y\begin{pmatrix}
 -1 \\
 1 \\
0
\end{pmatrix}" +"z\begin{pmatrix}
 -1 \\
 0\\
1
\end{pmatrix}"


rref of "\begin{pmatrix}
 -1 & -1 \\
 1 & 0\\
0&1
\end{pmatrix}" ="\begin{pmatrix}
 1& 0 \\
 0 & 1\\
0&0
\end{pmatrix}"


Basis are ["\begin{pmatrix}
 -1 \\
 1\\
0
\end{pmatrix}" , "\begin{pmatrix}
 -1 \\
 0\\
1
\end{pmatrix}" ]



Part 2

Number of linearly independent columns in "W" is 2

"\therefore" The dimension of W =2